OpenStudy (anonymous):

What is the area of a circle with a diameter of 16 maters? use 22/7 for pi.

OpenStudy (anonymous):

22/7*8^2

OpenStudy (anonymous):

Why would you use 22/7 for pi?!? Just use pi!

OpenStudy (anonymous):

RIGHT!!??!!??!!??!!??!!

OpenStudy (anonymous):

22/7 is approximate of pi

OpenStudy (anonymous):

so wats the answer.

OpenStudy (anonymous):

The area of a circle is \[\pi \times r^2\], And the radius is half the diameter (thus, 8) \[22/7 \times 8^2 = ~201.142848\]

OpenStudy (anonymous):

There is no reason to approximate it. Certainly not with a ratio..

OpenStudy (anonymous):

\[r = 8 \implies Area = \pi (8^2) = 64\pi \]

OpenStudy (anonymous):

HELP!!!!

OpenStudy (anonymous):

Just tell me the answer

OpenStudy (anonymous):

If you don't want it approximated, that's 1408/7, or 201 1/7